OpenStudy (anonymous):

an equation of the line containing the given point and parallel to the given line (3,6); x+7y=2

OpenStudy (aroub):

parallel=same slope -7y=-x+2 y=1/7x-2/7 y=mx+b , m=1/7 , (3,6) 6=1/7(3)+b 6=3/7+b -3/7+42/7 =b 39/7=b y=1/7x+39/7

OpenStudy (anonymous):

how this one same problem with just different numbers, )-6.5)2x=3y+7

OpenStudy (aroub):

okay!

OpenStudy (aroub):

y=mx+b , m=2/3 , (-6,5) 5=2/3(-6)+b 5=-4+b 9=b y=2/3x+9